# From Pen to Prompt: How Creative Writers Integrate AI into their Writing Practice

### v2

@@ 4.1.2. When are writers cautious about using AI?
Writers want to retain a feeling of ownership and control. Writers locate ownership in different aspects of the writing process, leading to varying strategies for maintaining authorial control. For some, ownership centers on the initial creative vision. P2 explains, “That story premise was my idea. The words—AI wrote them, but I feel ownership of the idea.” Importantly, writers adjust their AI use when it begins to detract from their enjoyment of the process. For instance, when P2 offloaded too much to the AI, using it to create the main plot in a new genre, the sense of ownership degraded; P2 decided to reduce AI use in their next project. While writers often want AI to fit their style, when it mimics their style well, writers may actually feel less ownership. For P7: “it doesn’t feel like 100% me… 30% to 35% could easily be attributed to the AI, and I think part of that is because the AI is so good at mimicking the style and extending it.”
@@
Writers want to preserve their enjoyment of the craft. Respondents view writing as a site of engagement, not automation. They avoided using AI to generate text or ideas without first putting in their own effort or ensuring that significant editing and rewriting were involved. This commitment to the craft of writing spurs a reluctance to having AI do the work for them. As P2 explains, “my personal satisfaction is lower when the AI does most of the work…”. P10 further states, “If you weren’t a good writer writing stuff by hand, you’re not going to be a good writer with AI. It’s just going to extrapolate your foolishness”. Understanding the writing process, genre expectations, and being able to distinguish and produce quality writing are important to writers. P6 notes that, for AI-assisted works, “you get to judge it on the quality. So if an AI wrote it… this guy was not involved in the process at all. He just clicked the button… it’s a story, but is it a decent story?” Even when AI provides suggestions or generates text, writers believe they must still “put in the effort for it to come out with magnificent results” (P6), applying rounds of revision and decisions to align with their vision.
@@ 4.2. How Are Writers Incorporating AI into their Writing Process?
4.2.1. How do writers determine their relationship with AI?
@@
While AI might perform similar tasks across different relationships (both assistant and collaborator roles can derive plot ideas), how writers frame these relationships shapes their overall approach to working with AI. Below, we summarize writers’ relationships with AI and organize them based on their perceived power dynamics, from those where writers maintain explicit authority (boss/assistant, director/actor) to more collaborative exchanges (fellow writer, editor) to inspirational (muse), to mechanical framings (tool). We provide more example quotes for every relationship expressed by participants in Appendix D.
Several participants described AI in subservient roles. As an assistant, AI is performing tasks given by the writer, who is the decision maker and boss. For example, P12 sees AI as “a very good assistant …kind of like a supportive yes man.” Similarly, as an actor, the writer is the director, guiding the AI to perform tasks or behave in certain ways that fit their needs and the vision they have in their heads. P6 says “I think more my role in using this is kind of like a director, and he is the actor. So I give it what it needs, then I see what happens, and then, if I like it, I keep it going.” AI output is molded according to the writer’s creative vision, much like an actor following a director’s cues. When AI is the assistant or actor, writers maintain tight control and desire high quality output.
In contrast, collaborative relationships akin to a fellow writer show greater receptiveness to AI’s creative contributions, often emerging as writers develop better expectations for AI’s capabilities. Writers describe AI as contributing ideas, writing text, and giving feedback in the capacity of another writer. These perceptions range from seeing AI as an active co-writer, “bringing its own things to the table” (P16) and helping to shape the text alongside the writer, through ideas, prose, or feedback in an iterative manner, to more passive roles such as junior writers, with less experience and skill, needing significant input and editing from the writer to be useful. Also mentioned was a ghostwriter, whose role is to help the author get their ideas down in their voice, contributing substantially less to the creative direction. Editor roles represent a middle ground where writers maintain creative control but are open to collaborative feedback. Writers use AI to identify strengths and weaknesses, restructure text, and give suggestions. P13 uses AI to simulate “the viewpoint of an editor for a writer before my essay gets published.
@@
4.2.2. How do writers integrate AI into their writing process?
Writers integrate AI into their workflows in ways that reflect their creative values, goals, and relationships with AI. We summarize the integration strategies writers deploy, such as how writers prompt, evaluate and incorporate text, in Figure 1 and in more detail in Figure 2. Overall, these strategies are not just technical processes but ways writers reconcile tensions, make choices, and experiment. A full list of all the tasks we uncovered for which writers use AI can be found in Appendix E.
Writers constantly make decisions around how to use AI in their workflow. Writers continuously make decisions about AI, from the initial setup (e.g., which platform to use) to individual writing moments (e.g., creating dialogue). Beginning with the decisions around setting up their writing environment, writers noted that different AIs have distinct characteristics: some are better for prose generation, others more constrained and reliable for research, while some are “crazy” and better for creative ideation. For example, P1 uses ChatGPT for research and only uses Sudowrite for writing when stuck.
Decision moments (Figure 1) are key to understanding how writers integrate AI into the creative process at various levels of granularity, from individual words to entire plot lines. During writing sessions, participants explained the moments where a writer might turn to AI. At any of these calls to AI, writers consider whether to use AI based on their vision and goals at each stage. When writers have a clear vision, they often choose to write directly to maintain their intended message. When their vision is less clear, writers may use AI to explore or clarify ideas while still maintaining editorial control over the final direction. These decision points allow writers to adjust AI use based on their values and the current role AI is taking on. At any stage, writers may also shift from a focus on writing to revision, using AI to assess and refine their text with actions such as summarizing, improving structure and clarity, or aligning with genre expectations. Writers might also turn to AI for enrichment, such as doing research or asking questions to deepen character motivations.
@@ 4.2.2. How do writers integrate AI into their writing process?
Evaluation is where writers exert considerable control. Writers evaluate AI output from multiple dimensions: alignment with their voice and vision, logical consistency, genre expectations, quality standards, and utility. Evaluation benefits from intuition built on experience writing, as P14 notes in evaluating alignment with style: “I can immediately evaluate a sentence as whether or not that’s a sentence I have the ability to write… I’m a good writer to write with AI, because I’ve already done my internship, my apprenticeship, and my years in the trenches.” If outputs don’t meet standards, writers iterate by adjusting prompts, modifying context, or providing additional information—until the writer receives something usable, and proceeds to incorporate the text, which often involves more edits on the part of the writer, and a closer inspection: “I have read every single word and I am putting my stamp of approval that this can go out the door.” (P14).
Unexpected generations can still be useful. Even when AI produces something writers weren’t aiming for, it can still trigger a “spark” (Gero et al., 2022)—an unexpected moment of inspiration that pushes the writer forward. As AI generates more material, the chance of something catching the writer’s attention increases. For P1: “just having something bad is useful, just to see how it flows.” The process of generating and rejecting options also helps writers clarify their own visions, where even “bad” outputs serve as valuable anti-patterns: showing writers what they don’t want helps them understand what they do want. As P10 notes, “Even poor quality outputs are still useful for sparking ideas.”
@@ 4.3. The Interplay of Values, Relationships, and Integration
The interactions between writers’ values, their relationships with AI, and their integration strategies serve to explain observed patterns in writer behavior, where each influences and reinforces the others (Figure 3). Values set the foundation of what writers find most important in the writing process, the core beliefs and principles that guide behavior. Writers choose when, why, and how to use AI in order to preserve and align with their values, from rewriting AI text to maintain authentic voice to creating experimental prompts for more creative exploration. Yet these strategies can also shift how writers experience their values over time and continued use. Writers can discover their initial boundaries (such as “writing every word myself” to preserve ownership) can evolve into more nuanced approaches as they use AI in new ways and still retain the feeling of ownership. On the other hand, when a writers’ use of AI conflicts with their values, writers will adjust their strategies, reinforcing the importance of their boundaries.
Values, which are individual to each writer and their literary genres, also strongly influence how they first approach AI and conceive of its role. For example, those prioritizing ownership often begin by viewing AI as a tool, while writers from literary traditions more open to referencing and remixing more readily embrace AI as a muse or creative collaborator. However, these relationships can evolve to challenge writers’ conceptions of their values, reconsidering what concepts like ownership and craftsmanship mean in practice, leading to further refinement of values.
How writers view their relationships with AI, whether as tool, collaborator, or muse, shapes their initial strategies for integration. Those who see AI as an assistant develop more task-specific workflows while those seeing AI as a muse might experiment more freely. The ways in which writers use AI in turn influence how writers understand AI’s role. As writers discover new and effective approaches or as more sophisticated AI models arise, their conception of AI’s role can expand. For example, many writers used ChatGPT and saw AI as an assistant due to its overly helpful tone and predictable responses, until they tried AI platforms for creative writing with fewer guardrails, which they then saw as more of a collaborator. Similarly, as writers run into challenges, their conceptions of AI as a limited assistant or tool can be reinforced.
@@ 5.1. Writers’ AI Workflows
The writers in our study demonstrate ways of working with AI that are far from automating the process, instead developing new workflows and expertise in areas such as prompting, evaluation, and creative experimentation. These new workflows suggest that rather than automating creative writing, AI may be creating new dimensions of writerly expertise. Writer workflows vary significantly across genres: poets often approach AI as a material generator to be reshaped, while fiction writers focus more on AI as productivity support. Personal essayists were most restrictive in not using AI to generate text, typically using AI only for research and feedback. Rather than converging on a single model of AI use, we may see continued divergence as different writing communities develop practices aligned with their values and traditions. This observation carries implications for new AI-infused writing tools: while scaffolded support may accelerate common tasks, it might also stymie more creative and genre-specific workflows.
@@ 5.3. Market Pressures, Productivity, and Creativity
The tension between productivity and creativity in writing is not new (Boice, 1993; Kachelmeier et al., 2008). Professional writers, particularly in genre fiction where publishing speed impacts income, described feeling pressure to maintain competitive publishing speeds. This suggests a potential future where AI adoption may become necessary for market participation, rather than truly optional. Some writers turn to AI due to health challenges to level the playing field, raising questions about escalating productivity expectations. Our participants expressed ethical concerns such as environmental impact, models’ training data source, and writer compensation, yet many balanced these pragmatically with productivity expectations. Although all participants hoped for a future where AI would support human creativity, some participants predicted a future where AI-generated commercial writing and human craft writing would become increasingly separate. Supporting sustainable creative careers may require new frameworks for valuing and compensating creative work that account for both human craft and technological assistance.
@@ 5.4. Transparency of New Processes
Additionally, broader societal stigma around AI-generated content make it even harder for writers to share their processes. Despite how writers would ideally like to give attribution, whether they choose to disclose AI becomes a separate issue, due to stigma and perceptions of AI-assisted writing. Unfortunately, stigma was even internalized, as P7 (who uses AI to help overcome carpal tunnel) said: “I do feel like an imposter for using AI in the sense that maybe I’m cheating a little bit, or maybe I’m not working as hard as I could be.”
Design and research direction: Supporting visibility of AI integration to share new processes. Rather than focusing on detecting or marking AI use, new tools should help writers document and optionally share their creative processes. Future interfaces could build in rich process documentation that enables writers to track decision points and creative choices in ways meaningful to them, serving both as personal reflection and potential knowledge sharing. These systems might support different levels of process sharing, from personal records to writing groups to public disclosure. The goal should not be surveillance, but supporting writers in understanding their own practice while enabling voluntary knowledge sharing within creative communities. This could help reduce stigma by moving discourse beyond binary AI or human debates toward more nuanced understanding of how writers maintain creative authenticity while using new tools. Over time, communities in different genres could develop new norms for AI attribution that extend from existing ones. For instance, P15 shared they come from a poetic literary tradition with an “unwritten code of ethics about referencing lines… It feels like poetry is very much a kind of a communal effort… So do I feel like [ownership] should be claimed?…it really depends on how the poem came about, and how heavily it leans on AI, and whether or not the knowing that AI was part of the process benefits the reading of the poem.” By making processes more transparent through voluntary sharing rather than policing, communities could develop best practices naturally through example and exposure.

### v4

@@ 4.1.2. When are writers cautious about using AI in their writing?
Writers want to retain a feeling of ownership and control. Writers locate ownership in different aspects of the writing process, leading to varying strategies for maintaining authorial control. For some, ownership centers on the initial creative vision. P2 explains, “That story premise was my idea. The words—AI wrote them, but I feel ownership of the idea.” Importantly, writers adjust their AI use when it begins to detract from their enjoyment of the process. For instance, P2 offloaded too much to the AI in trying to speed up the writing of a book by using AI to create the main plot in a new genre and their sense of ownership degraded; P2 then decided to reduce AI use in their next project. While writers often want AI to fit their style, when it mimics their style well, writers may actually feel less ownership. For P7: “it doesn’t feel like 100% me… 30% to 35% could easily be attributed to the AI, and I think part of that is because the AI is so good at mimicking the style and extending it.”
@@
Writers want to preserve their enjoyment of the craft. Respondents view writing as a site of engagement, not automation. They avoided using AI to generate text or ideas without first putting in their own effort or ensuring that significant editing and rewriting were involved. This commitment to the craft of writing spurs a reluctance to having AI do the work for them. As P2 explains, “my personal satisfaction is lower when the AI does most of the work…” P10 further states, “If you weren’t a good writer writing stuff by hand, you’re not going to be a good writer with AI. It’s just going to extrapolate your foolishness”. Understanding the writing process, genre expectations, and being able to distinguish and produce quality writing are important to writers. P6 notes that, for AI-assisted works, “you get to judge it on the quality. So if an AI wrote it… this guy was not involved in the process at all. He just clicked the button… it’s a story, but is it a decent story?” Even when AI provides suggestions or generates text, writers believe they must still “put in the effort for it to come out with magnificent results” (P6), applying rounds of revision and decisions to align with their vision.
@@ 4.2. How Are Writers Incorporating AI into their Writing Process?
4.2.1. How do writers determine their relationship with AI in practice?
@@
While AI might perform similar tasks across different relationships (both assistant and collaborator roles can derive plot ideas), how writers frame these relationships shapes their overall approach to working with AI. Below, we summarize writers’ relationships with AI and organize them based on their perceived power dynamics, from those where writers maintain explicit authority (boss/assistant, director/actor) to more collaborative exchanges (fellow writer, editor) to inspirational (muse), to mechanical framings (tool). These roles were named by the writers, with their characterizations further informed by both the interviews and observed writing sessions. We provide more example quotes for every relationship expressed by participants in Appendix F.
Several participants described AI in subservient roles. As an assistant, AI is performing tasks given by the writer (decision maker and boss). For example, P12 sees AI as “a very good assistant …kind of like a supportive yes man.” Similarly, as an actor, the writer is the director, guiding the AI to perform tasks or behave in certain ways that fit their needs and the vision they have in their heads. P6 says “I think more my role in using this is kind of like a director, and he is the actor. So I give it what it needs, then I see what happens, and then, if I like it, I keep it going.” AI output is molded according to the writer’s creative vision, much like an actor following a director’s cues. When AI is the assistant or actor, writers maintain tight control and desire high quality output.
In contrast, collaborative relationships akin to a fellow writer show greater receptiveness to AI’s creative contributions, often emerging as writers develop better expectations for AI’s capabilities. Writers describe AI as contributing ideas, writing text, and giving feedback in the capacity of another writer. These perceptions range from seeing AI as an active collaborator, “bringing its own things to the table” (P16) and helping to shape the text alongside the writer, through ideas, prose, or feedback in an iterative manner, to more passive roles such as junior writers, with less experience and skill, needing significant input and editing from the writer to be useful. Also mentioned was a ghostwriter, whose role is to help the author get their ideas down in their voice, contributing substantially less to the creative direction. Editor roles represent a middle ground where writers maintain creative control but are open to collaborative feedback. Writers use AI to identify strengths and weaknesses, restructure text, and give suggestions. P13 uses AI to simulate “the viewpoint of an editor for a writer before my essay gets published.
@@
4.2.2. How are writers integrating AI into their writing process?
Writers integrate AI into their workflows in ways that reflect their creative values, goals, and relationships with AI. We summarize the integration strategies writers deploy, such as how writers prompt, evaluate and incorporate text, in Figure 1 and in more detail in Figure 2. Overall, these strategies are not just technical processes but ways writers reconcile tensions, make choices, and experiment. A full list of all the tasks we uncovered for which writers use AI can be found in Appendix G.
Writers constantly make decisions around how to use AI in their workflow. Writers continuously make decisions about AI, from the initial setup (e.g., which platform to use) to individual writing moments (e.g., creating dialogue). These decisions are shaped by prior experience and evolving familiarity with different AI, as writers develop a sense of which platforms and prompting strategies best serve their needs. Beginning with the decisions around setting up their writing environment, writers noted that different AIs have distinct characteristics: some are better for prose generation, others more constrained and reliable for research, while some are “crazy” and better for creative ideation. For example, P1 uses ChatGPT for research and only uses Sudowrite for writing when stuck. Seventeen out of 18 of the writers had used or started with ChatGPT as the most accessible entry point. As their needs for creativity, stylistic control, and consistency over longform texts grew, many began experimenting with more sophisticated prompts or branched into platforms with more specialized features for creative writing.
Decision moments (Figure 1) are key to understanding how writers integrate AI into the creative process at various levels of granularity, from individual words to entire plot lines. During writing sessions, participants explained the moments where a writer might turn to AI. At any of these calls to AI, writers consider whether to use AI based on their vision and goals at each stage. When writers have a clear vision, they often choose to write directly to maintain their intended message. When their vision is less clear, writers may use AI to explore or clarify ideas while still maintaining editorial control over the final direction. For example, P13 did not have a specific vision for their song but rather started with a general topic, using AI to surface relevant themes, characters, and settings. A phrase from this exploration struck their inspiration and ultimately became the song’s title. These decision points allow writers to adjust AI use based on their values and the current role AI is taking on. At any stage, writers may also shift from a focus on writing to revision, using AI to assess and refine their text with actions such as summarizing, improving structure and clarity, or aligning with genre expectations. Writers might also turn to AI for enrichment, such as doing research or asking questions to deepen character motivations.
@@ 4.2.2. How are writers integrating AI into their writing process?
Evaluation is where writers exert considerable control. Writers evaluate AI output from multiple dimensions: alignment with their voice and vision, logical consistency, genre expectations, quality standards, and utility. Evaluation benefits from intuition built on experience in writing, as P14 notes in evaluating alignment with style: “I can immediately evaluate a sentence as whether or not that’s a sentence I have the ability to write… I’m a good writer to write with AI, because I’ve already done my internship, my apprenticeship, and my years in the trenches.” If outputs don’t meet standards, writers iterate by adjusting prompts, modifying context, or providing additional information—until the writer receives something usable, and proceeds to incorporate the text, which often involves more edits on the part of the writer, and a closer inspection: “I have read every single word and I am putting my stamp of approval that this can go out the door” (P14).
Unexpected generations can still be useful. Even when AI produces something writers weren’t aiming for, it can still trigger a “spark” (Gero et al., 2022)—an unexpected moment of inspiration that pushes the writer forward. As AI generates more material, the chance of something catching the writer’s attention increases. For P1: “just having something bad is useful, just to see how it flows.” The process of generating and rejecting options also helps writers clarify their own visions, where even “bad” outputs serve as valuable anti-patterns: showing writers what they don’t want helps them understand what they do want. As P10 notes, “even poor quality outputs are still useful for sparking ideas.” Sparks often emerged unexpectedly: prompts that generated more text material to work with or surfaced a wider range of ideas were more likely to yield sparks (asking for themes (P13), putting characters in dialogue with each other (P1), or reverse-prompting, where AI asked the writer questions back to reflect on their writing (P4)). For the genre of poetry especially, P16 explained “all the mistakes, or the things that feel unnatural, or unexpected, they’re immediately so much richer than something that feels familiar: anything that sounds slightly odd, like it sticks, kind of rings in the air. You want to work with it, and it suggests new ideas as well.”
@@ 4.3. The Interplay of Values, Relationships, and Integration
The interactions between writers’ values, their relationships with AI, and their integration strategies serve to explain observed patterns in writer behavior, where each influences and reinforces the others (Figure 3).
Values and Integration Strategies. Writers choose when, why, and how to use AI in order to preserve and align with their values, from rewriting AI text to maintain authentic voice to creating experimental prompts for more creative exploration. Yet these strategies can also shift how writers experience their values over time and continued use. Writers can discover their initial boundaries (such as “writing every word myself” to preserve ownership) can evolve into more nuanced approaches as they use AI in new ways: “I think that it’s only reinforced my idea of originality and authentic self… you can absolutely be an authentic artist as long as it’s your work and it’s your interpretation” (P13). On the other hand, when a writers’ use of AI conflicts with their values, writers will adjust their strategies, reinforcing the importance of their boundaries, as in the case of P2 reverting their AI strategy to preserve feeling of ownership.
Values and Relationships with AI. Values, which are individual to each writer and their literary genres, also strongly influence how they first approach AI and conceive of its role. For example, those prioritizing ownership often begin by viewing AI as a tool, while writers from literary traditions more open to referencing and remixing more readily embrace AI as a muse or creative collaborator. However, these relationships can evolve to challenge writers’ conceptions of their values, reconsidering what concepts like ownership and craftsmanship mean in practice, leading to further refinement of values. P5, for example, began by seeing AI as a tool, but came to describe it as a “ghostly co-writer” with its own direction, prompting them to feel like an “imposter,” and ultimately reframe ownership as something shared rather than singular.
Relationships with AI and Integrations Strategies. How writers view their relationships with AI, whether as tool, collaborator, or muse, shapes their initial strategies for integration. Those who see AI as an assistant develop more task-specific workflows while those seeing AI as a muse might experiment more freely. For instance, P8’s view of AI as a human companion directly influenced their decision not to use any AI-generated text directly: “I like to treat [AI] as a human being… and if I were to include any line from ChatGPT, I just feel like I can’t credit it, so I won’t use it.” The ways in which writers use AI in turn influence how writers understand AI’s role. As writers discover new and effective approaches or as more sophisticated AI models arise, their conception of AI’s role can expand. For example, many writers used ChatGPT and saw AI as an assistant due to its overly helpful tone and predictable responses, until they tried AI platforms for creative writing with fewer guardrails, which they then saw as more of a collaborator. Similarly, as writers run into challenges, their conceptions of AI as a limited assistant or tool can be reinforced.
@@ 5.1. Writers’ AI Workflows
The writers in our study demonstrate ways of working with AI that are far from automating the process, instead developing new workflows and expertise in areas such as prompting, evaluation, and creative experimentation. These new workflows suggest that rather than automating creative writing, AI may be creating new dimensions of writerly expertise. Writer workflows vary significantly across genres: poets often approach AI as a material generator to be reshaped, being more comfortable in experimenting with prompts. Fiction writers focus more on AI as productivity support, for tasks like ideation and prose generation. Personal essayists were most restrictive in not using AI to generate text, typically using AI only for research and feedback. Rather than converging on a single model of AI use, we may see continued divergence as different writing communities develop practices aligned with their values and traditions. This observation carries implications for new AI-infused writing tools: while scaffolded support may accelerate common tasks, it might also stymie more creative and genre-specific workflows.
@@ 5.3. Market Pressures, Productivity, and Creativity
The tension between productivity and creativity in writing is not new (Boice, 1993; Kachelmeier et al., 2008). Professional writers, particularly in genre fiction where publishing speed impacts income, described feeling pressure to maintain competitive publishing speeds with P10 stating “we get paid by the word, the more books I put out, the more money I make.” This suggests a potential future where AI adoption may become necessary for competitive market participation in some genres, rather than truly optional especially as AI has been found to reshape job content as it amplifies productivity (Eloundou et al., 2024; Autor, 2022). Some writers turn to AI due to health challenges to level the playing field, raising questions about escalating productivity expectations. Our participants expressed ethical concerns such as environmental impact, models’ training data source, and writer compensation, yet many balanced these pragmatically with productivity expectations. Although all participants hoped for a future where AI would support human creativity, some participants predicted a future where AI-generated commercial writing and human writing would become increasingly separate: “I think it’s going to become like a craft art, if that makes sense. Like, for example, this shirt was woven on a machine loom…” (P14). Supporting sustainable creative careers will require new frameworks for valuing and compensating creative work that account for both human craft and technological assistance.
@@ 5.4. Transparency of New Processes
Additionally, broader societal stigma around AI-generated content make it even harder for writers to share their processes. Despite how writers would ideally like to give attribution, whether they choose to disclose AI becomes a separate issue, due to stigma and perceptions of AI-assisted writing. Unfortunately, stigma was even internalized, as P5 said: “I do feel like an imposter for using AI in the sense that maybe I’m cheating a little bit, or maybe I’m not working as hard as I could be.”
Design and research direction: Supporting visibility of AI integration to share new processes. Rather than focusing on detecting or marking AI use, new tools should help writers document and optionally share their creative processes. Future interfaces could build in rich process documentation that enables writers to track decision points and creative choices in ways meaningful to them, serving both as personal reflection and potential knowledge sharing. These systems might support different levels of process sharing, from personal records to writing groups to public disclosure. The goal should not be surveillance, but supporting writers in understanding their own practice while enabling voluntary knowledge sharing within creative communities. This could help reduce stigma by moving discourse beyond binary AI or human debates toward more nuanced understanding of how writers maintain creative authenticity while using new tools. Over time, communities in different genres could develop new norms for AI attribution that extend from existing ones. For instance, P15 shared they come from a poetic literary tradition with an “unwritten code of ethics about referencing lines… It feels like poetry is very much a kind of a communal effort… So do I feel like [ownership] should be claimed?…it really depends on how the poem came about, and how heavily it leans on AI, and whether or not the knowing that AI was part of the process benefits the reading of the poem.” By making processes more transparent through voluntary sharing rather than policing, future research could explore how such visibility helps communities develop best practices by revealing diverse workflows, fostering discussion, and modeling examples of responsible AI use.